OpenStudy (anonymous):

Fan and medal Which number is halfway between 3/4 and 3/5 ?

OpenStudy (anonymous):

@aaronq

OpenStudy (anonymous):

@jdoe0001

OpenStudy (dan815):

okay to find a number between

OpenStudy (dan815):

you have to look at the average of both numbers

OpenStudy (dan815):

do you know how to find average of 2 numbers?

OpenStudy (anonymous):

no do you add them?

OpenStudy (dan815):

yeah you add the numbers and divide by the total number of numbers you added so you add 2 numbers 3/4,3/5 you divided by 2 after you add them (3/4 + 3/5) / 2 = ?

OpenStudy (dan815):

\[\frac{\frac{3}{4} +\frac{3}{5} }{2}\]

OpenStudy (anonymous):

when i added them i got 27/20 is the first part right?

OpenStudy (anonymous):

Is it 27/40

OpenStudy (jdoe0001):

now as dan815 said , divide them by 2 \(\bf \cfrac{\frac{3}{4} +\frac{3}{5} }{2}\implies \cfrac{\frac{27}{20}}{2}\implies \cfrac{\frac{27}{20}}{\frac{2}{1}}\implies \cfrac{27}{20}\cdot \cfrac{1}{2}\)

OpenStudy (jdoe0001):

recall that \(\bf \cfrac{\frac{a}{b}}{\frac{c}{{\color{blue}{ d}}}}\implies \cfrac{a}{b}\cdot \cfrac{{\color{blue}{ d}}}{c}\)

OpenStudy (anonymous):

would it be 2 7/10

OpenStudy (jdoe0001):

\(\bf \cfrac{\frac{27}{20}}{\frac{2}{1}}\implies \cfrac{27}{20}\cdot \cfrac{1}{2}\implies \cfrac{27\cdot 1}{20\cdot 2}\)

OpenStudy (anonymous):

Oh so it would be 27/40

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

Thank you

OpenStudy (jdoe0001):

yw